Prepare 240 g of 3.5% sulfate ointment using a 2% ointment and a 4% ointment.
Answer by Alan3354(69443)

Prepare 240 g of 3.5% sulfate ointment using a 2% ointment and a 4% ointment.
-------------
t = 2%
f = 4%
-----
t + f = 240
2t + 4f = 3.5*240